OpenStudy (anonymous):

Solve this system of equations using the addition method. .5m + .7n = 6.4 .2m - .7n = .6 @mathstudent55

OpenStudy (whpalmer4):

Add the two equations together, like you're adding some multi-digit numbers on paper: \[.5m + .7n = 6.4\]\[.2m-.7n = .6\]---------------- \[.7m + 0n = 7.0\] Can you solve for \(m\) now?

OpenStudy (anonymous):

no

OpenStudy (whpalmer4):

why not?

OpenStudy (whpalmer4):

\[.7m = 7.0\]

OpenStudy (whpalmer4):

Divide both sides by the number in front of the letter you are solving for. \[\frac{.7m}{.7} = \frac{7.0}{.7}\]\[m=\]

OpenStudy (anonymous):

m=10 right

OpenStudy (whpalmer4):

exactly! that wasn't so hard, was it?

OpenStudy (whpalmer4):

now we have to find the value of \(n\). we can plug \(m = 10\) into either one of the equations and find \(n\)

OpenStudy (anonymous):

but tht wasnt the right answer

OpenStudy (anonymous):

but still i dont get ut

OpenStudy (whpalmer4):

what do you mean it wasn't the right answer? what do you think the right answer is?

OpenStudy (anonymous):

idk

OpenStudy (whpalmer4):

why are you saying it isn't the right answer?

OpenStudy (anonymous):

i check it and and its not

OpenStudy (whpalmer4):

You checked it wrong :-) Let's continue on to find \(n\) if \(m=10\) \[.5m + .7n = 6.4\] (first equation) \[.5(10) + .7n = 6.4\]\[5+.7n=6.4\]Subtract 5 from both sides\[.7n=1.4\]divide both sides by .7\[n = 2\] Now we check our work: \[.5(10)+.7(2) = 5 + 1.4 = 6.4\checkmark\]Must also check any other equations!\[.2(10)-.7(2) = 2-1.4=0.6\checkmark\]

OpenStudy (anonymous):

so whats the answer

OpenStudy (whpalmer4):

We need to check the solutions in all of the equations in the system because it is possible to find a set of numbers which will satisfy some of the equations but not all of them. For example, m=12.8, n = 0 satisfies the first equation: .5(12.8)+.7(0) = 6.4 but it does not satisfy the second one: .2(12.8)-.7(0) = 2.56

OpenStudy (whpalmer4):

Read my posts again, you'll see the answer.

OpenStudy (anonymous):

2.56

OpenStudy (whpalmer4):

no... Come on, read through them from the beginning of the thread. I came up with a value for m which you insisted was wrong. Then I found the value for n, and I checked them both. Look at the equations which have checkmarks and you'll see the numbers.

OpenStudy (whpalmer4):

This is a system of equations with two variables (m, n), so we will have 2 numbers as "the answer" (m=, n=).